OpenStudy (dragontears1128):

can someone help me please (i want to know what you got for the answer) The annual Gross Domestic Product (GDP) of a country is the value of all of the goods and services produced in the country during a year. During the period 1985-1999, the Gross Domestic Product of the United States grew about 3.2% per year, measured in 1996 dollars. In 1985, the GDP was $577 billion. I what year did/or will the GDP equal $3.4 trillion?

OpenStudy (dragontears1128):

i got 2088 and other numbers but its wrong

OpenStudy (dragontears1128):

@Slenderman

OpenStudy (slenderman):

there is a discrepancy in the q since GDP is in $, one would think 1 billion = 10^9, 1 trillion = 10^12 but the GDP for the base year 1985 has been given as $5717 billion = $5.717 trillion it can't GROW from that to $ 2 trillion anyway, let's try the british convention, so we have 2*10^18 = 5.717*10^15 (1.0032)^n 1.0032^n = 2*10^3/5.717 = 349.8338 n*ln1.0032 = ln 349.8338 n = ln 349.8338 / ln 1.0032 = 1833.38 ans: 1834th year from 1985 ----- p.s: ---- the q makes no sense. i'm pretty sure the figures have been garbled

OpenStudy (slenderman):

@TheSmartOne

OpenStudy (dragontears1128):

like every time i anwer it it says its wrong

OpenStudy (dragontears1128):

so i messaged my teacher